OpenStudy (anonymous):

The parabola y=x^2+3x-18 crosses the x-axis at

OpenStudy (saifoo.khan):

Solve the quadratic!

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

its different how would i start my problem

OpenStudy (saifoo.khan):

Okay. Let me teach you in 2 minutes.

OpenStudy (saifoo.khan):

Consider this equation: ax^2 + bx + c What we do, we look for TWO terms whom when we MULTPLY, we get the value of a*c And when we ADD those same terms, we get the value of b! Read this statement 3 times, try to understand. Tell me when we agree till here.

OpenStudy (anonymous):

okay im good from here wats next

OpenStudy (saifoo.khan):

Alright! So in your case: x^2+3x-18 a = 1, b = 3 and c = -18. Agree?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Okay. So: a*c = 1*-18 = -18 AND b = 3 So now we have find TWO terms (negative or positive), which when we ADD we get 3 and when we MULTIPLY we get -18. Agree?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Perfect. Now start searching those terms. +6 and -3. Check: 6*-3 = -18 (a*c) 6 + (-3) = +3 (b)

OpenStudy (anonymous):

Saifoo.Khan outstanding, Factoring gives (x+6)(x-3) =0. The product of 6*-3 = -18, their sum is +3. The roots are the negative of those factors.

OpenStudy (anonymous):

can it be 3 and -6 as well

OpenStudy (saifoo.khan):

Yes. Since we get 2 roots, we rewrite as: x^2 +6x - 3x - 18 =0 then make pairs, x(x+6) -3(x+6) = 0 Finally: (x - 3)(x+6) = 0

OpenStudy (saifoo.khan):

@jmays14 : No! It can't be! 6 - 3 = 3 -6 + 3 = -3

OpenStudy (anonymous):

so that means the parabola doesnt cross the x axis

OpenStudy (saifoo.khan):

It does! at x = +3 and x = -6

OpenStudy (anonymous):

okay these are the answer choices that i have but -6 and 3 isnt one (A) 3 and -6 (B) 2 and -6 (C) 3 and -9 (D) 2 and -9 (E) The parabola does not cross the x-axis.

OpenStudy (anonymous):

i mean 6,-3

OpenStudy (saifoo.khan):

It's A!

OpenStudy (anonymous):

oh okay can u hlep me with 1 more please

OpenStudy (saifoo.khan):

Umm. Okay. Post it on left.